OpenStudy (anonymous):

If the relation y = (1/2)x + 5 has the range {4, 6, 9}, then what is the corresponding domain? A. {-2, 2, 8} B. {-2, 2, 13} C. {3, 7, 13} D. {-7, -3, 3}

OpenStudy (anonymous):

4=(1/2)x+5 x=-2 6=(1/2)x+5 x=2 9=(1/2)x+5 x=8 answer: a